OpenStudy (anonymous):

factor the expression and use fundamental identities to simplify. sin^2xcsc^2x-sin^2x

OpenStudy (lgbasallote):

hmm will knowing \[\large \csc^2 x = \frac{1}{\sin^2 x}\] help?

OpenStudy (anonymous):

how is the soution cos^2x

OpenStudy (lgbasallote):

\[\Large \sin^2 x \times \frac{1}{\sin^2 x} - \sin^2 x\] do you agree with this?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

and what's \[\sin^2 x \times \frac{1}{\sin^2 x}\]

OpenStudy (anonymous):

i thought we could cancel them :o

OpenStudy (lgbasallote):

yup...so what would it be?

OpenStudy (anonymous):

1? but the answer is cos^2x :o

OpenStudy (lgbasallote):

lol we'll get there :p

OpenStudy (lgbasallote):

\[(\sin^2 x \times \frac{1}{\sin^2 x}) - \sin^2x\] \[\implies (1) - \sin^2 x\] agree?

OpenStudy (anonymous):

ywa

OpenStudy (lgbasallote):

so what's \(1 - \sin^2 x?\)

OpenStudy (anonymous):

im guessing cos^2x LOL but i dont see it in my identities

OpenStudy (lgbasallote):

\[\sin^2 x + \cos^2 x = 1\] do you see that in your identities?

OpenStudy (anonymous):

ohhh justi kidding! Pythagoream identities

OpenStudy (lgbasallote):

that was pythagorean identity? lol...fancy terms =)))

OpenStudy (anonymous):

yeah ahaha